OpenStudy (anonymous):

How do you find the equation of a parabola with zero's (-3,0) and (4,0)?

OpenStudy (anonymous):

kinda work backward from the steps you would take to solve for the zeros... if x=-3 is a zero then x+3 could be a factor because x+3=0------>x=-3 same for the next one, x-4=0 so multiply together to get the equation y=(x-4)(x+3)